OpenStudy (anonymous):

integral of 1/x^2 (e^2/x)

OpenStudy (agent0smith):

\[\LARGE \int\limits \frac{ 1 }{ x^2 } e ^{\frac{ 2 }{x }} dx\] Let u = 2/x

OpenStudy (gorv):

well 1/x=t can u solve further???

OpenStudy (anonymous):

e^2/x is in the denominator

OpenStudy (agent0smith):

Then there should have been parentheses around it :/ \[\LARGE \int\limits\limits \frac{ 1 }{ x^2 e ^{\frac{ 2 }{x }} } dx\] Still, let \[\large u = \frac{ 2 }{ x }\]

OpenStudy (gorv):